the area of a rectangle is 96 square centimeters. if the width of the rectangle is 4 centimeter less than its length, what are the dimension of the rectangle
Answer by mananth(16946)

length= x
width = x-4
...
area = L*W
Area = x(x-4)=96
x^2-4x=96
x^2-4x-96=0
x^2-12x+8x-96=0
x(x-12)+8(x-12)=0
*x-12)(x+8)=0
So x= 12 cms===>length
width = 12-4=8 cm